Instructions: Solve each system by the substitution method.
What is the solution for 5x+2y=-1
2x-y=-13

Answer by kayeith(17)

5x + 2y = -1
2x - y = -13
a) - y = - 2x - 13
-1 (-y = - 2x - 13)
y = 2x + 13
b) 5x + 2 (2x + 13) = -1
5x + 4x + 26 = -1
9x = -1 - 26
9x = -27
x = -27 / 9
x = -3
c) y = 2x + 13
y = 2 (-3) + 13
y = -6 + 13
y = 7